solve the equation. round any solutions to the nearest fourth decimal place.
4ln(2x)+ 3ln(x^5)=9

Answer by stanbon(75887)

4ln(2x)+ 3ln(x^5)=9
---
ln(2x)^4 + ln(x)^15 = 9
---
ln[16x^5*x^19] = 9
---
16x^24 = e^9
---
x^24 = [e^9/16]
---
x = 1.2962
===============
Cheers,
Stan H.
=============